OpenStudy (anonymous):

Can someone Plzzzzzzzzzzzzzzzzzzz HELP!!!!!! Match Term Definition three x equals four A) two x equals three x divided by twelve equals four B) six x equals eight x divided by eight equals five C) x divided by four equals twelve four x equals six D) x divided by four equals ten

OpenStudy (anonymous):

someone plz plz help me !!!!!

OpenStudy (anonymous):

I will give a medal and fan you

jigglypuff314 (jigglypuff314):

well we can start by turning it all into math speak XD 3x = 4 A. 2x = 3 x/12 = 4 B. 6x = 8 x/8 = 5 C. x/4 = 12 4x = 6 D. x/4 = 10

jigglypuff314 (jigglypuff314):

then you would have to solve for x for each :) to do that, you want to get x all alone on one side so like for 3x = 4 you would divide both sides by 3 so that x = 4/3 and for something like x/12 = 4 you would multiply both sides by 12 so that x = 4 times 12 = 48

OpenStudy (anonymous):

thank you so much @jiggleypuff314

jigglypuff314 (jigglypuff314):

you're welcome :) you know how to do it now?

jigglypuff314 (jigglypuff314):

you solve for x for all of them then match the answers from the left to the ones on the right :) so like x/12 = 4 matches with C

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so like this @jigglepuff314 A) 3x=4 C)x/12=4 B)x/8=4 D)4x=6

jigglypuff314 (jigglypuff314):

mmm try again?

jigglypuff314 (jigglypuff314):

3x = 4 -> x = 4/3 A. 2x = 3 -> x = 3/2 x/12 = 4 -> x = 48 B. 6x = 8 -> x = 4/3 x/8 = 5 -> x = 40 C. x/4 = 12 -> x = 48 4x = 6 -> x = 3/2 D. x/4 = 10 -> x = 40 so B = 3x=4 C = x/12=4 D = x/8=4 A = 4x=6

OpenStudy (anonymous):

thank you

jigglypuff314 (jigglypuff314):

glad I could help :)